For the elemental sulfur combustion : (1/8)S + O2 --> SO2

Find:

a) \DeltaSo298K

b) \DeltaHo298 K

c) KP at 298 K and at 773 K. (Assume that the temperature dependence of the values of ?S
Answer #1

Delta S = delta S ( SO2) - [1/8 deltaS ( sulphur) + delta S (O2)]

= 248.11 - [1/8(167.74) + 205.03] = 22.11 J/K

Delta H = delta H ( SO2) - [1/8 delta H( sulphur) + delta H (O2)]

Delta H = -296.83 - [0+0] = -296.83

c) delta G = deltaH-TdeltaS = -296.83 - 773 X 0.02211 = -313.92

dG = -2.303RT logKp

logKp = -313.92 / 8.314 X 773 = -0.05

solve for Kp